OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of 2 and contains the point (3, 7).

OpenStudy (anonymous):

The slope intercept equation is given by : y-y1=m(x-x1), where m is the slope. W

OpenStudy (anonymous):

where m is the slope and the coordinates correspond to x1 and y1

OpenStudy (anonymous):

Plugging in the given values give us: y-7=2(x-3). Distribute the 2 on the right side and we get: y-7=2x-6. Then we add 7 to both sides of the equation: y=2x+1

OpenStudy (anonymous):

slope= 2 and coordinates (3,7) so y=mx+b 7=2(3)+b 7=6 6- 7=1 y=2x+1